# Finding Most Shattering Minimum Vertex Cuts of Polylogarithmic Size in Near-Linear Time

### v1

@@ Abstract
We show the first near-linear time randomized algorithms for listing all minimum vertex cuts of polylogarithmic size that separate the graph into at least three connected components (also known as shredders) and for finding the most shattering one, i.e., the one maximizing the number of connected components. Our algorithms break the quadratic time bound by Cheriyan and Thurimella (STOC’96) for both problems that has been unimproved for more than two decades. Our work also removes an important bottleneck to near-linear time algorithms for the vertex connectivity augmentation problem (Jordan ’95) and finding an even-length cycle in a directed graph, a problem shown to be equivalent to many other fundamental problems (Vazirani and Yannakakis ’90, Robertson et al. ’99). Note that it is necessary to list only minimum vertex cuts that separate the graph into at least three components because there can be an exponential number of minimum vertex cuts in general.
@@ 1 Introduction
Given an $n$-vertex, $m$-edge, undirected graph $G$, a minimum vertex cut is a set of vertices whose removal disconnects $G$ of smallest possible size. The vertex connectivity of $G$ is the size of any minimum vertex cut. The problem of efficiently computing vertex connectivity and finding a corresponding minimum vertex cut has been extensively studied for more than half a century [Kle69, Pod73, ET75, Eve75, Gal80, EH84, Mat87, BDD+82, LLW88, CT91, NI92, CR94, Hen97, HRG00, Gab06, CGK14]. Let $k$ denote the vertex connectivity of $G$. Recently, an $\mathcal{O}(k^{2}m)$-time algorithm was shown in [FNS+20], which is near-linear when $k=\mathcal{O}(\text{polylog}(n))$. Then, an almost-linear $m^{1+o(1)}$-time algorithm for the general case was finally discovered [LNP+21, CKL+22]. In this paper, we show new algorithms for two closely related problems:
@@
In the latter problem, the restriction to at least three components is natural for polynomial-time algorithms. This is because there are at most $n$ many minimum vertex cuts whose removal results in at least three components [Jor99], while the total number of minimum vertex cuts can be exponential or, more specifically, at least $2^{k}(n/k)^{2}$ [PY21]. [^1] Both problems are specific to vertex cuts; recall that every minimum edge cut always separates a graph into two components. Before we discuss the literature on both problems, let us introduce some terminology. We say that a vertex set $S$ is a separator if $G\setminus S$ is not connected and a shredder if $G\setminus S$ has at least three components. An $s$-separator ($s$-shredder) is a separator (shredder) of size $s$. In other words, the second problem is to list all $k$-shredders. [^2] For example, in a tree, vertices with degree at least three are 1-shredders. In the complete bipartite graph $K_{k,k}$ with $k\geq 3$, both the left and right halves of the bipartition are $k$-shredders.
@@
The first efficient algorithm for computing most shattering min-cuts (Problem 1) was by Cheriyan and Thurimella [CT99]. They showed an algorithm with $\mathcal{O}(k^{2}n^{2}+k^{3}n\sqrt{n})$ running time and used it as a subroutine to speed up Jordán’s approximation algorithm mentioned above to $\mathcal{O}(\min(k,\sqrt{n})k^{2}n^{2}+kn^{2}\log n)$ time. [^3] Years later, Végh [Vég11] gave the first exact polynomial-time algorithm that runs in $\mathcal{O}(kn^{7})$ time using a different approach. A more general problem was also studied: given an arbitrary parameter $k^{\prime}$, we must determine the minimum number of edges required to increase the vertex connectivity of $G$ from $k$ to $k^{\prime}$. Although polynomial-time algorithms are still not known for this problem, Jackson and Jordán [JJ05] showed a fixed-parameter-tractable algorithm that runs in $\mathcal{O}(n^{6}+c_{k}n^{3})$ time, where $c_{k}$ is an exponential value that depends only on $k$. The parameter $b(G)$ is also important in both their algorithm and analysis.
@@
Hegde’s motivation behind his algorithm is a surprising connection to the problem of finding an even-length cycle in a directed graph. Vazirani and Yannakakis [VY89] showed that this problem is equivalent to many other fundamental questions, e.g., checking if a bipartite graph has a Pfaffian orientation (see [Kas63, Kas67]) and more. Hegde observed via the work of [RST99] that the bottleneck to speed up the even-length cycle algorithm beyond $\mathcal{O}(n^{2})$ time is determining whether a 4-connected bipartite graph contains a 4-shredder. However, Hegde’s algorithm only works for $3$-shredders, and the fastest known algorithm even for $4$-shredders is by [CT99], which still requires $\Omega(n^{2})$ time.
The algorithm by [CT99] for both Problems 1 and 2 has remained state-of-the-art for more than 20 years. Moreover, its approach inherently requires quadratic time because it makes $\Omega(k^{2}+n)$ max flow calls. Unfortunately, this quadratic running time has become a bottleneck for the related tasks mentioned above. Naturally, one may wonder: do subquadratic algorithms exist for both problems?
@@
Our paper is organized as follows. Section˜2 describes the bottleneck of the algorithm presented in [CT99] and our high-level casework approach to resolving this bottleneck. We also solve the simpler case and explain why the remaining case requires more sophisticated methods. Section˜3 provides the preliminaries needed throughout the paper. Section˜4 explains the core algorithm in [CT99] and restates important definitions and terminology. Sections˜5, 6 and 7 are devoted to solving the hard case of this problem introduced in Section˜2. Section˜8 summarizes the previous sections and shows the algorithm for listing all $k$-shredders. Section˜9 provides the algorithm for finding the most shattering min-cut. Lastly, Section˜10 discusses the overall results and implications for open problems.
@@ 2 Technical Overview
Let $G=(V,E)$ be an $n$-vertex $m$-edge undirected graph with vertex connectivity $k$. Cheriyan and Thurimella developed a deterministic algorithm called $\texttt{All-$k$-shredders}(\cdot)$ that takes $G$ as input and lists all $k$-shredders of $G$ in $\mathcal{O}(knm+k^{2}\sqrt{n}m)$ time. They improved this bound by using the sparsification routine developed in [NI92] as a preprocessing step. Specifically, there exists an algorithm that takes $G$ as input and produces an edge subgraph $G^{\prime}$ on $\mathcal{O}(kn)$ edges such that all $k$-shredders of $G$ are $k$-shredders of $G^{\prime}$ and vice versa. The algorithm runs in $\mathcal{O}(m)$ time. Using this preprocessing step, they obtained the bound for listing all $k$-shredders in $\mathcal{O}(m)+\mathcal{O}(k(kn)n+k^{2}\sqrt{n}(kn))=\mathcal{O}(k^{2}n^{2}+k^{3}n\sqrt{n})$ time.
In this paper, we resolve a bottleneck of $\texttt{All-$k$-shredders}(\cdot)$, improving the time complexity of listing all $k$-shredders from $\mathcal{O}(knm+k^{2}\sqrt{n}m)$ to $\mathcal{O}(k^{4}m\log^{4}n)$. Using the same sparsification routine, our algorithm runs in $\mathcal{O}(m+k^{5}n\log^{4}n)$ time.
@@ 2.1 The Bottleneck
A key subroutine of $\texttt{All-$k$-shredders}(\cdot)$ is a subroutine called $\texttt{Shredders}(\cdot,\cdot)$ that takes a pair of vertices $(x,y)$ as input and lists all $k$-shredders that separate $x$ and $y$. This subroutine takes $\mathcal{O}(m)$ time plus the time to compute a flow of size $k$, which is at most $\mathcal{O}(km)$ time. The idea behind $\texttt{All-$k$-shredders}(\cdot)$ is to call $\texttt{Shredders}(\cdot,\cdot)$ multiple times to list all $k$-shredders. $\texttt{All-$k$-shredders}(\cdot)$ works as follows. Let $Y$ be an arbitrary set of $k+1$ vertices. Any $k$-shredder $S$ will either separate a pair of vertices in $Y$, or separate a component containing $Y\setminus S$ from the rest of the graph.
@@ 2.3 Listing Balanced $k$-Shredders via Edge Sampling
Listing all balanced $k$-shredders turns out to be straightforward via random edge sampling. Let $S$ be a $k$-shredder that admits partition $(C,\mathcal{R})$. Suppose that $S$ is balanced. The idea is to sample two edges $(x,x^{\prime})$ and $(y,y^{\prime})$ such that $x$ and $y$ are in different components of $G\setminus S$. Then, $\texttt{Shredders}(x,y)$ will list $S$. Intuitively, this approach works because we know that each component of $G\setminus S$ cannot be too large, as $S$ is a balanced $k$-shredder. More precisely, by sampling $\mathcal{O}(k\log n)$ pairs of edges, with high probability, at least one pair of edges will hit different connected components because the biggest component $C$ has volume at most mm/km-m/k. The formal statement is given below. Since its proof is standard, we defer it to Appendix˜A.
@@ 2.4 Listing Unbalanced $k$-Shredders via Local Flow
What is most important about this result is that we have constructed an algorithm that can identify $k$-shredders on a local scale, i.e., proportional to an input volume parameter $\nu$. The idea behind Lemma˜2.5 is to modify $\texttt{Shredders}(\cdot,\cdot)$ using recent advancements in local flow algorithms.
A major caveat concerns the set $U$. We can think of $U$ as an unverified set. Essentially, because the algorithm in Lemma˜2.5 is time-limited by a volume parameter $\nu$, we may discover a set $U$ but not have enough time to verify whether $U$ is a $k$-shredder.
@@ 3 Preliminaries
This paper concerns finite, undirected, and unweighted graphs with vertex connectivity $k$. We use standard graph-theoretic definitions. Let $G=(V,E)$ be a graph with vertex connectivity $k$. For a vertex subset $S\subseteq V$, we use $G\setminus S$ to denote the subgraph of $G$ induced by removing all vertices in $S$. A connected component (component for short) of a graph refers to any maximally-connected subgraph, or the vertex set of such a subgraph. Suppose that $S$ is a $k$-shredder of $G$. The largest component of $G\setminus S$ is the component with the greatest volume, where volume is defined in Definition˜2.1. We break ties arbitrarily. For convenience, we will say that a $k$-shredder $S$ admits partition $(C,\mathcal{R})$ to signify that $C$ is the largest component of $G\setminus S$ and $\mathcal{R}$ is the set of remaining components. We will also write $x\in\mathcal{R}$ to indicate a vertex in a component of $\mathcal{R}$.
@@ 4 Cheriyan and Thurimella’s Algorithm
At a high level, $\texttt{Shredders}(\cdot,\cdot)$ works as follows. Let $x$ and $y$ be two vertices in a $k$-vertex-connected graph $G$. Firstly, we use a flow algorithm to obtain a set $\Pi$ of $k$ openly-disjoint simple paths from $x$ to $y$. Observe that any $k$-shredder $S$ of $G$ separating $x$ and $y$ must contain exactly one vertex from each path of $\Pi$. More crucially, at least one component of $G\setminus S$ must also be a component of $G\setminus\Pi$. That is, for some component $Q$ of $G\setminus\Pi$, we have $N(Q)=S$. This is the main property that $\texttt{Shredders}(\cdot,\cdot)$ exploits. It lists potential $k$-shredders by finding components $Q$ of $G\setminus\Pi$ such that $|N(Q)|=k$ and $N(Q)$ consists of exactly one vertex from each path in $\Pi$ (see Figure˜1).
Figure 1: Let $\Pi$ be the set of three openly-disjoint simple paths from $x$ to $y$. A call to $\texttt{Shredders}(x,y)$ will identify $N(Q_{1})=N(Q_{2})$ as potential 33-shredders.
@@
If $\Gamma$ is an edge, we define $\mathrm{vol}(\Gamma)=1$. Otherwise, $\mathrm{vol}(\Gamma)$ is defined according to Definition˜2.1. For a set of bridges $\mathcal{B}$, we define $\mathrm{vol}(\mathcal{B})=\sum_{\Gamma\in\mathcal{B}}\mathrm{vol}(\Gamma)$.
@@
Definition 4.6 ($\preceq,\succeq$).
@@
For a visual description of Definition˜4.7, see figure Figure˜2. It will also be useful to notice that the statement $\Gamma_{1}$ straddles $\Gamma_{2}$ is equivalent to $\Gamma_{1}\not\preceq\Gamma_{2}$ and $\Gamma_{1}\not\succeq\Gamma_{2}$. Furthermore, Definitions˜4.6 and 4.7 are still well-defined even for candidates (i.e. we can use the $\preceq,\succeq$ operators to compare bridges to candidates and candidates to candidates).
@@
Note that Lemma˜4.9 generalizes $\texttt{Shredders}(\cdot,\cdot)$ in the sense that it does not require that all paths in $\Pi$ were between a pair of vertices $(x,y)$. This will be particularly useful in the later sections. Because the proofs of Lemmas˜4.8 and 4.9 directly follow from the ideas from [CT99], we defer their proofs along with the proof of ˜4.1 to Appendix˜B.
@@ 5 Locally Listing $k$-Shredders or an Unverified Set
[…]
@@
To prove Lemma˜2.5, we first state some structural properties of shredders when captured by $(x,\nu,\Pi)$ in Section˜5.1. Then, we describe the algorithm of in Section˜5.2 and analyze it in Section˜5.3.
@@ 5.1 Structures of Local Shredders
The following lemmas translate the structural properties of shredders from Section˜4 to the setting when they are captured by the tuple $(x,\nu,\Pi)$. They will be needed in the analysis of our algorithm.
@@
The following statement is the local analog to the statement discussed below ˜4.1.
@@
By ˜5.1, $\Pi$ cannot contain a vertex in any component of $\mathcal{R}\setminus\{Q_{x}\}$, as such a path would necessarily use more than one vertex of $S$. Thus, for an arbitrary component $Q\in\mathcal{R}\setminus\{Q_{x}\}$, we have $\Pi\cap Q=\varnothing$. It follows that $Q\subseteq Q^{\prime}$ for some component $Q^{\prime}$ of $G\setminus\Pi$.
@@
Next, we prove one direction of the “if and only if” statement of Lemma˜4.8 in the local setting. Another direction requires knowing how our algorithm constructs candidate $k$-shredders and will be proved later in Lemma˜5.9.

### v2

@@ Abstract
We show the first near-linear time randomized algorithms for listing all minimum vertex cuts of polylogarithmic size that separate the graph into at least three connected components (also known as shredders) and for finding the most shattering one, i.e., the one maximizing the number of connected components. Our algorithms break the quadratic time bound by Cheriyan and Thurimella (STOC’96) for both problems that has stood for more than two decades. Our work also removes a bottleneck to near-linear time algorithms for the vertex connectivity augmentation problem (Jordan ’95). Note that it is necessary to list only minimum vertex cuts that separate the graph into at least three components because there can be an exponential number of minimum vertex cuts in general.
@@ 1 Introduction
Given an $n$-vertex, $m$-edge, undirected graph $G$, a minimum vertex cut is a set of vertices whose removal disconnects $G$ of smallest possible size. The vertex connectivity of $G$ is the size of any minimum vertex cut. The problem of efficiently computing vertex connectivity and finding a corresponding minimum vertex cut has been extensively studied for more than half a century [Kle69, Pod73, ET75, Eve75, Gal80, EH84, Mat87, BDD+82, LLW88, CT91, NI92, CR94, Hen97, HRG00, Gab06, CGK14]. Let $k$ denote the vertex connectivity of $G$. Recently, an $\mathcal{O}(k^{2}m)$ time algorithm was shown in [FNS+20], which is near-linear when $k=\mathcal{O}(\text{polylog}(n))$. Then, an almost-linear $m^{1+o(1)}$ time algorithm for the general case was finally discovered [LNP+21, CKL+22]. In this paper, we show new algorithms for two closely related problems:
@@
In the latter problem, the restriction to at least three components is natural for polynomial time algorithms. This is because there are at most $n$ many minimum vertex cuts whose removal results in at least three components [Jor99], while the total number of minimum vertex cuts can be exponential or, more specifically, at least $2^{k}(n/k)^{2}$ [PY21]. [^1] Both problems are specific to vertex cuts; recall that every minimum edge cut always separates a graph into two components. Before we discuss the literature on both problems, let us introduce some terminology. We say that a vertex set $S$ is a separator if $G\setminus S$ is not connected and a shredder if $G\setminus S$ has at least three components. An $s$-separator ($s$-shredder) is a separator (shredder) of size $s$. In other words, the second problem is to list all $k$-shredders. [^2] For example, in a tree, vertices with degree at least three are 1-shredders. In the complete bipartite graph $K_{k,k}$ with $k\geq 3$, both the left and right halves of the bipartition are $k$-shredders.
@@
The first efficient algorithm for computing most shattering min-cuts (Problem 1) was by Cheriyan and Thurimella [CT99]. They showed an algorithm with $\mathcal{O}(k^{2}n^{2}+k^{3}n^{3/2})$ running time and used it as a subroutine to speed up Jordán’s approximation algorithm mentioned above to $\mathcal{O}(\min(k,\sqrt{n})k^{2}n^{2}+kn^{2}\log n)$ time. [^3] Years later, Végh [Vég11] gave the first exact polynomial time algorithm that runs in $\mathcal{O}(kn^{7})$ time using a different approach. A more general problem was also studied: given an arbitrary parameter $k^{\prime}$, we must determine the minimum number of edges required to increase the vertex connectivity of $G$ from $k$ to $k^{\prime}$. Although polynomial time algorithms are still not known for this problem, Jackson and Jordán [JJ05] showed a fixed-parameter-tractable algorithm that runs in $\mathcal{O}(n^{6}+c_{k}n^{3})$ time, where $c_{k}$ is an exponential value that depends only on $k$. The parameter $b(G)$ is also important in both their algorithm and analysis.
@@
Hegde’s motivation behind his algorithm is a surprising connection to the problem of finding an even-length cycle in a directed graph. Vazirani and Yannakakis [VY89] showed that this problem is equivalent to many other fundamental questions, e.g., checking if a bipartite graph has a Pfaffian orientation (see [Kas63, Kas67]) and more. Via the work of [RST99], Hegde observed that to speed up the algorithm for finding Pfaffian orientations in bipartite graphs beyond $\mathcal{O}(n^{2})$ time, one needs to quickly determine whether a special type of bipartite graph contains a 4-shredder. [^4] It turns out that this special type of bipartite graph (defined as a brace by [RST99]) is 3-connected when it contains at least five vertices, but not necessarily 4-connected. To truly remove this bottleneck, one must at the very least determine whether a 3-connected brace contains a 4-shredder. Therefore, our algorithm does not remove this bottleneck. In the previous version of this paper, we misunderstood Hegde’s statement and incorrectly claimed that we removed this bottleneck. Up until now, the fastest known algorithm for listing 4-shredders in 4-connected graphs is by [CT99], which still requires $\Omega(n^{2})$ time.
The algorithm by [CT99] for both Problems 1 and 2 has remained state-of-the-art for more than 20 years. Moreover, its approach inherently requires quadratic time because it makes $\Omega(k^{2}+n)$ max flow calls. Unfortunately, its quadratic runtime has become a bottleneck for the related tasks mentioned above. Naturally, one may wonder: do subquadratic algorithms exist for both problems?
@@
Our paper is organized as follows. Section 2 describes the bottleneck of the algorithm presented in [CT99] and our high-level casework approach to resolving this bottleneck. We also solve the simpler case and explain why the remaining case requires more sophisticated methods. Section 3 provides the preliminaries needed throughout the paper. Section 4 explains the core algorithm in [CT99] and restates important definitions and terminology. Sections 5, 6 and 7 are devoted to solving the hard case of this problem introduced in Section 2. Section 8 summarizes the previous sections and shows the algorithm for listing all $k$-shredders. Section 9 provides the algorithm for finding the most shattering min-cut. Lastly, Section 10 discusses the overall results and implications for open problems.
@@ 2 Technical Overview
[…]
@@ 2.1 The Bottleneck
[…]
@@ 2.3 Listing Balanced $k$-Shredders via Edge Sampling
Listing all balanced $k$-shredders turns out to be straightforward via random edge sampling. Let $S$ be a $k$-shredder that admits partition $(C,\mathcal{R})$. Suppose that $S$ is balanced. The idea is to sample two edges $(x,x^{\prime})$ and $(y,y^{\prime})$ such that $x$ and $y$ are in different components of $G\setminus S$. Then, $\texttt{Shredders}(x,y)$ will list $S$. Intuitively, this approach works because we know that each component of $G\setminus S$ cannot be too large, as $S$ is a balanced $k$-shredder. More precisely, by sampling $\mathcal{O}(k\log n)$ pairs of edges, with high probability, at least one pair of edges will hit different connected components because the biggest component $C$ has volume at most $m-m/k$. The formal statement is given below. Since its proof is standard, we defer it to Appendix A.
@@ 2.4 Listing Unbalanced $k$-Shredders via Local Flow
What is most important about this result is that we have constructed an algorithm that can identify $k$-shredders on a local scale, i.e., proportional to an input volume parameter $\nu$. The idea behind Lemma 2.5 is to modify $\texttt{Shredders}(\cdot,\cdot)$ using recent advancements in local flow algorithms.
A major caveat concerns the set $U$. We can think of $U$ as an unverified set. Essentially, because the algorithm in Lemma 2.5 is time-limited by a volume parameter $\nu$, we may discover a set $U$ but not have enough time to verify whether $U$ is a $k$-shredder.
@@ 3 Preliminaries
This paper concerns finite, undirected, and unweighted graphs with vertex connectivity $k$. We use standard graph-theoretic definitions. Let $G=(V,E)$ be a graph with vertex connectivity $k$. For a vertex subset $S\subseteq V$, we use $G\setminus S$ to denote the subgraph of $G$ induced by removing all vertices in $S$. A connected component (component for short) of a graph refers to any maximally-connected subgraph, or the vertex set of such a subgraph. Suppose that $S$ is a $k$-shredder of $G$. The largest component of $G\setminus S$ is the component with the greatest volume, where volume is defined in Definition 2.1. We break ties arbitrarily. For convenience, we will say that a $k$-shredder $S$ admits partition $(C,\mathcal{R})$ to signify that $C$ is the largest component of $G\setminus S$ and $\mathcal{R}$ is the set of remaining components. We will also write $x\in\mathcal{R}$ to indicate a vertex in a component of $\mathcal{R}$.
@@ 4 Cheriyan and Thurimella’s Algorithm
At a high level, $\texttt{Shredders}(\cdot,\cdot)$ works as follows. Let $x$ and $y$ be two vertices in a $k$-vertex-connected graph $G$. Firstly, we use a flow algorithm to obtain a set $\Pi$ of $k$ openly-disjoint simple paths from $x$ to $y$. Observe that any $k$-shredder $S$ of $G$ separating $x$ and $y$ must contain exactly one vertex from each path of $\Pi$. More crucially, at least one component of $G\setminus S$ must also be a component of $G\setminus\Pi$. That is, for some component $Q$ of $G\setminus\Pi$, we have $N(Q)=S$. This is the main property that $\texttt{Shredders}(\cdot,\cdot)$ exploits. It lists potential $k$-shredders by finding components $Q$ of $G\setminus\Pi$ such that $|N(Q)|=k$ and $N(Q)$ consists of exactly one vertex from each path in $\Pi$ (see Figure 1).
Figure 1: Let $\Pi$ be the set of three openly-disjoint simple paths from $x$ to $y$. A call to $\texttt{Shredders}(x,y)$ will identify $N(Q_{1})=N(Q_{2})$ as potential $3$-shredders.
@@
If $\Gamma$ is an edge, we define $\mathrm{vol}(\Gamma)=1$. Otherwise, $\mathrm{vol}(\Gamma)$ is defined according to Definition 2.1. For a set of bridges $\mathcal{B}$, we define $\mathrm{vol}(\mathcal{B})=\sum_{\Gamma\in\mathcal{B}}\mathrm{vol}(\Gamma)$.
@@
Definition 4.6 ($\preceq$,$\succeq$ , ⪰).
@@
For a visual description of Definition 4.7, see figure Figure 2. It will also be useful to notice that the statement $\Gamma_{1}$ straddles $\Gamma_{2}$ is equivalent to $\Gamma_{1}\not\preceq\Gamma_{2}$ and $\Gamma_{1}\not\succeq\Gamma_{2}$. Furthermore, Definitions 4.6 and 4.7 are still well-defined even for candidates (i.e. we can use the $\preceq$,$\succeq$ , ⪰ operators to compare bridges to candidates and candidates to candidates).
@@
Note that Lemma 4.9 generalizes $\texttt{Shredders}(\cdot,\cdot)$ in the sense that it does not require that all paths in $\Pi$ were between a pair of vertices $(x,y)$. This will be particularly useful in the later sections. Because the proofs of Lemmas 4.8 and 4.9 directly follow from the ideas from [CT99], we defer their proofs along with the proof of 4.1 to Appendix B.
@@ 5 Locally Listing $k$-Shredders or an Unverified Set
Suppose we have obtained a tuple $(x,\nu,\Pi)$ that captures a $k$-shredder $S$. Recall Definition 2.4 for the definition of capture. The method for constructing $(x,\nu,\Pi)$ that captures $S$ is via random sampling and is deferred to Section 7. In this section, we show a local algorithm that, assuming $S$ is captured by $(x,\nu,\Pi)$, puts $S$ in the returned list $\cal{L}$ or reports $S$ as an unverified set $U$. How we handle the unverified set will be explained in Section 6.
@@
To prove Lemma 2.5, we first state some structural properties of shredders when captured by $(x,\nu,\Pi)$ in Section 5.1. Then, we describe the algorithm of in Section 5.2 and analyze it in Section 5.3.
@@ 5.1 Structures of Local Shredders
The following lemmas translate the structural properties of shredders from Section 4 to the setting when they are captured by the tuple $(x,\nu,\Pi)$. They will be needed in the analysis of our algorithm.
@@
The following statement is the local analog to the statement discussed below 4.1.
@@
By 5.1, $\Pi$ cannot contain a vertex in any component of $\mathcal{R}\setminus\{Q_{x}\}$, as such a path would necessarily use more than one vertex of $S$. Thus, for an arbitrary component $Q\in\mathcal{R}\setminus\{Q_{x}\}$, we have $\Pi\cap Q=\varnothing$. It follows that $Q\subseteq Q^{\prime}$ for some component $Q^{\prime}$ of $G\setminus\Pi$.
@@
Next, we prove one direction of the “if and only if” statement of Lemma 4.8 in the local setting. Another direction requires knowing how our algorithm constructs candidate $k$-shredders and will be proved later in Lemma 5.9.